Hello , help me to solve [ Question 2 ]

les are Lo egl O tythe total number sfmia state of the assembly n averayt ocupa 2 k: YE! 3 2 E 0.123 1.15 스터-173 Wh: 27 I 27enele
Answer #1

e Ua CE amd maxi mum enent avel e level GE 3 0 The 320 2 4 eThiefm So, the ento m The enena lelel se e U t and maximun 32 0he mo e a the change im enlo 41 - 23lnA+lnB=ln(AB)

lnA-lnB=ln(\frac{A}{B})